Charles lives in Detroit and runs a business that sells pianos. In an average year, he receives $716,000 from selling pianos. Of this sales revenue, he must pay the manufacturer a wholesale cost of $416,000; he also pays wages and utility bills totaling $274,000. He owns his showroom; if he chooses to rent it out, he will receive $7,000 in rent per year. Assume that the value of this showroom does not depreciate over the year. Also, if Charles does not operate this piano business, he can work as a financial advisor and receive an annual salary of $22,000 with no additional monetary costs. No other costs are incurred in running this piano business. Identify each of Charless costs in the following table as either an implicit cost or an explicit cost of selling pianos. Implicit Cost Explicit Cost The rental income Charles could receive if he chose to rent out his showroom The wages and utility bills that Charles pays The salary Charles could earn if he worked as a financial advisor The wholesale cost for the pianos that Charles pays the manufacturer Complete the following table by determining Charless accounting and economic profit of his piano business. Profit (Dollars) Accounting Profit Economic Profit Alternatively, the economic profit he would earn as a financial advisor would be If Charless goal is to maximize his economic profit, he stay in the piano business True or false: Charles is not earning a normal profit be not earning a normal proft beca should is negative. should not False True
Answer #1

Implicit cost is an opportunity cost that is what firms give up in order to do another work.

An explicit cost is an actual cost which incurs during the current work

The rental income Charles could receive if he chooses to rent out his showroom- Implicit cost.

The wage and utility bills that Charles pays- Explicit cost.

The salary Charles could earn if he worked as a financial advisor- Implicit cost

The wholesale cost of the pianos that Charles pays the manufacturer- Explicit cost

Accounting profit of his piano buisiness= 716000-(416000+274000)= $26000

Economic profit= Accounting profit- Implicit cost= 26000-7000= $19000

Economic profit if he would earn as a financial advisor= 26000-22000= $4000

If Charles's goal is to maximize his economic profit he should stay in piano business. As the economic profit from piano business is more than that of working as a financial advisor

Charles is not earning a normal profit because his earning is more than both explicit and implicit cost. The given statement is false.